Find z1/n; for n=3, z = 1-i in C (, the Argand Plane).

1
Expert's answer
2021-12-06T16:25:58-0500

The polar form of "1−i=\sqrt{2}(\cos(-\dfrac{\pi}{4}+i\sin(-\dfrac{\pi}{4}))."

According to the De Moivre's Formula, all 3-th roots of a complex number 

"\sqrt{2}(\cos(-\dfrac{\pi}{4}+i\sin(-\dfrac{\pi}{4}))" are given by


"(\sqrt{2})^{1/3}(\cos(\dfrac{-\pi/4+2\pi k}{3})+i\sin(\dfrac{-\pi/4+2\pi k}{3})),""k=0,1,2"

"k=0:"


"(2)^{1/6}(\cos(\dfrac{-\pi/4+2\pi (0)}{3})+i\sin(\dfrac{-\pi/4+2\pi (0)}{3}))"

"=(2)^{1/6}(\cos(\dfrac{\pi}{12})-i\sin(\dfrac{\pi}{12}))"

"k=1:"


"(2)^{1/6}(\cos(\dfrac{-\pi/4+2\pi (1)}{3})+i\sin(\dfrac{-\pi/4+2\pi (1)}{3}))"

"=(2)^{1/6}(\cos(\dfrac{7\pi}{12})+i\sin(\dfrac{7\pi}{12}))"

"k=2:"


"(2)^{1/6}(\cos(\dfrac{-\pi/4+2\pi (2)}{3})+i\sin(\dfrac{-\pi/4+2\pi (2)}{3}))"

"=(2)^{1/6}(\cos(\dfrac{5\pi}{4})+i\sin(\dfrac{5\pi}{4}))"

"=(2)^{-1\/3}(1+i)"